OpenStudy (anonymous):

how do you calculate 4sin2(theta) without a calculator?

OpenStudy (anonymous):

what is the value of theta??

OpenStudy (anonymous):

u can use formula...sin2a=2sinacosa

OpenStudy (anonymous):

theta = pi/6 thanks!

OpenStudy (anonymous):

then ur question reduces to 8 sin theta * cos theta..

OpenStudy (anonymous):

sin (2*pi/6)=sin (pi/3)=sqrt(3)/2 4 sin (pi/3)=2 sqrt(3)

OpenStudy (anonymous):

how can i reduce sin4theta?

OpenStudy (anonymous):

sin4theta=2sin 2theta* cos 2 theta

OpenStudy (anonymous):

again\[\cos2 \theta =\sqrt{1-\sin^ 2 2\theta}\]

OpenStudy (anonymous):

thanks!